Use first principle method/definition method to obtain the derivative f(x) = 2 – x


1
Expert's answer
2021-09-12T23:51:31-0400
"f'(x)=\lim\limits_{h\to 0}\dfrac{f(x+h)-f(x)}{h}"

"=\lim\limits_{h\to 0}\dfrac{2-(x+h)-(2-x)}{h}"


"=\lim\limits_{h\to 0}\dfrac{-h}{h}=-2"

"f'(x)=(2-x)'=-1, x\in \R"